# mini-vec2vec: Scaling Universal Geometry Alignment with Linear Transformations

### v3

@@ 1 Introduction
Sentence embeddings capture rich semantic structure in high-dimensional vector spaces. They encode complex linguistic relationships and enable downstream tasks, such as semantic similarity and information retrieval. However, different architectures, training procedures, and even different models with the same architecture produce embedding spaces that cannot be compared directly, because they use a different coordinate system, a different “internal language”. The problem of aligning embedding spaces without parallel supervision – that is, without pairs of matching encodings of the same data – has emerged as a fundamental challenge with far-reaching implications. Successful unsupervised alignment has multiple implications across many domains: multilingual natural language processing (conneau2017word; zhang-etal-2017-adversarial), cross-modal alignment between text and image representations (maniparambil2024; schnaus2025), and privacy concerns (song2020informationleakage; text2vec).
@@
The recent vec2vec method (jha2025vec2vec) has demonstrated that unsupervised alignment is indeed possible using CycleGAN (zhu2017cyclegan). Their approach is theoretically grounded in a principle called the Universal Geometry or the Platonic Representation Hypothesis (huh2024platonic), which posits that well-trained representations, regardless of their specific architectures or training procedures, tend to converge toward geometrically similar spaces, even if their coordinate systems differ. This means that there is a way to rotate the coordinate systems so they “understand” each other – and vec2vec does precisely that; it identifies a rotation that makes the coordinate systems aligned. While vec2vec has proven effective, adversarial training frameworks like CycleGAN bring inherent challenges that limit their practical applicability. Training them is computationally intensive, requires careful hyperparameter tuning, and is subject to training instabilities like mode collapse, oscillatory behavior, or failure to converge (saxena2023gan_survey; goodfellow2017nips2016). Moreover, they typically demand substantial computational resources, including GPU acceleration and large amounts of training data – factors that may prohibit their utilization by most researchers and users.
Our point of departure is still Universal Geometry, but we take a different approach. We are motivated by the remarkable success of supervised alignment approaches, particularly neural network stitching methods (lenc2014stitching; bansal2021revisiting), which have demonstrated that simple affine transformations often suffice to align representations. But unfortunately, stitching assumes access to pairs of matching representations. Motivated by the Platonic Representation Hypothesis, we hypothesize that we can nonetheless uncover this linear alignment without even needing adversarial training. Specifically, we hypothesize that the similarity in the geometric structure – the relative arrangement of semantic clusters, the preservation of neighborhood structures, and the overall geometric organization of concepts – should provide sufficient signal for alignment without parallel supervision and without need for adversarial methods. To this end, we utilize the Relative Representations framework of moschella2022relative, which allows us to operate in a universal space agnostic of arbitrary coordinate systems. We then use it as a stepping stone to produce a noisy alignment that we refine with further steps. [^1] In Section 6, we discuss a variant that sidesteps the use of relative representations, but we tested it far less.
@@
mini-vec2vec is a simple and robust pipeline with competitive alignment performance offering several key advantages over adversarial approaches: computational efficiency (very fast and running on commodity CPU hardware), training stability (good performance and robustness to hyperparameters), interpretability (each step has a clear geometric interpretation), and sample efficiency (effective with substantially fewer training examples). Through comprehensive experiments, we demonstrate that this simple approach not only matches but often exceeds the performance of the more complex adversarial method while requiring orders of magnitude less computational resources.
@@ 2 Preliminaries
The hypothesis that neural networks converge to geometrically similar representations has been explored across multiple domains and scales. Early work by colah2015visualizing has revealed striking structural similarities in representations of different convolutional networks, [^2] Olah briefly refers to even earlier ideas (e.g., a concept called “meta-tSNE” by Hinton), but we failed to locate them. suggesting that the geometric organization of learned features might be more universal than previously assumed. The Platonic Representation Hypothesis, formally articulated by huh2024platonic, represents the most comprehensive theoretical framework for understanding Universal Geometry. The hypothesis posits that well-trained representations converge toward a shared geometric structure that reflects the underlying structure of the data domain. Having the same geometry implies that the distance between objects should be independent of the models themselves. For example, if one wants to compute the semantic similarity of “I love your dog” and “I love my cat” using cosine similarity between their embeddings, the answer should be roughly the same regardless of the underlying embedding model. Intuitively, this empirical observation stands to reason, as semantic similarity originates from the semantic meaning that comes from the data itself. Below, we discuss the framework of moschella2022relative that extends this observation into a powerful tool.
@@
moschella2022relative have operationalized the observations of colah2015visualizing and proposed to use relative representations as a mathematically principled framework for comparing latent spaces. While absolute coordinates vary arbitrarily across models, we know that relative information, such as similarity between two objects, is (approximately) agnostic to the choice of embedding model. They propose to use a shared set of anchor points that other data points will be compared to as a way to ground the position of data points in a model-agnostic way.
Denote the space of data points by $\mathcal{Z}$ (e.g., sentences, images, or any other modality) and the embedding function by $E:\mathcal{X}\rightarrow\mathbb{R}^{d}$. Given a set of anchor data points $\{\mathbf{a}_{i}\}_{i=1}^{k}\subseteq\mathcal{Z}$, the relative representation of $\mathbf{z}\in\mathcal{Z}$ under $E$ is:
$\mathbf{r}_{E}(\mathbf{z})=[f(E(\mathbf{z)},E(\mathbf{a}_{1}));f(E(\mathbf{z}),E(\mathbf{a}_{2}));\ldots;f(E(\mathbf{z}),E(\mathbf{a}_{k}))],$
where $f(\cdot,\cdot)$ is a similarity function (typically cosine similarity) and ; represents concatenation.
The power of relative representations is that, as shown by colah2015visualizing and moschella2022relative, for two encoders $E_{1},E_{2}$, the relative representations of the same data point is roughly the same, i.e., $\mathbf{r}_{E_{1}}(\mathbf{z})\approx\mathbf{r}_{E_{2}}(\mathbf{z})$. In other words, the function $\mathbf{r}_{E}:\mathcal{X}\rightarrow\mathbb{R}^{D}$ is itself an embedding that maps a data point into $k$-dimensional space in a way that is approximately independent of $E$ (for a sufficiently well-trained $E$) and only depends on the choice of anchor points.
3 Problem Formulation
We follow the setup laid out in vec2vec. We assume access to two pools of text embeddings, each from a different encoder, and our task is to align their embedding spaces without paired examples. Similar to vec2vec, we learn a parameterized mapping (neural network or linear transformation) rather than a one-to-one matching between data points. This approach offers several advantages:
  • Generalization: New embeddings can be translated without re-applying the algorithm
  • Robustness: Works when one-to-one point matching doesn’t exist, requiring only similar data distributions
  • Out-of-distribution capability: The parameterized nature can (potentially) enable generalization beyond the training distribution
The Universal Geometry hypothesis indicates that such a task can be possible: “good enough” models learn equivalent geometries – relative similarities between data points remain approximately the same regardless of the embedding model.
Let $\mathcal{X}_{A}=\{\mathbf{x}^{A}_{i}\}_{i=1}^{n_{A}}$ and $\mathcal{X}_{B}=\{\mathbf{x}^{B}_{j}\}_{j=1}^{n_{B}}$ denote two sets of $d$-dimensional embeddings, where $\mathbf{x}^{A}_{i},\mathbf{x}^{B}_{j}\in\mathbb{R}^{d}$. The sentences come from the same underlying distribution $\mathcal{D}$, but – unlike most other works – there is no overlap between the sentences themselves – only the distribution is shared. We organize these embeddings into matrices $\mathbf{X}_{A}\in\mathbb{R}^{n_{A}\times d}$ and $\mathbf{X}_{B}\in\mathbb{R}^{n_{B}\times d}$, where each row corresponds to a single embedding vector. The fundamental challenge is that while both spaces may contain similar geometric structures, they are related by an unknown transformation.
Our goal is to learn a transformation function $f:\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that $f(E_{A}(\mathbf{z}))\approx E_{B}(\mathbf{z})$ where $\mathbf{z}$ represents a datapoint $\mathbf{z}\in\mathcal{Z}$. More generally, the setup described in jha2025vec2vec allows for learning a pair of functions $f_{A},f_{B}$ sending embeddings to a third space, such that $f_{A}(E_{A}(\mathbf{z}))\approx f_{B}(E_{B}(\mathbf{z}))$. We will not take advantage of this flexibility in this paper. [^3] It also complicates the analysis of the results. For example, $f_{A}=f_{B}=\text{const}$ would be a trivial solution if we don’t impose additional constraints. Similarity then needs to be assessed more carefully.
@@ 4 Methodology
4.1 Motivation
In this work, we restrict our attention to linear transformations, i.e., $f(\mathbf{x})=\mathbf{W}\mathbf{x}$ where $\mathbf{W}\in\mathbb{R}^{d\times d}$. Moreover, we will constrain our search space to orthogonal transformations, i.e., matrices $\mathbf{W}$ that satisfy $\mathbf{W}\mathbf{W}^{T}=\mathbf{W}^{T}\mathbf{W}=\mathbf{I}$. This is motivated by the Universal Geometry hypothesis – we expect cosine similarities to be preserved across embedding spaces. Later on, we will relax this assumption and ultimately the matrix will be an exponentially-smoothed average of orthogonal matrices, enforcing orthogonality softly.
Our linear approach, mini-vec2vec, addresses key limitations of the current approach: GANs are notoriously unstable and hard to train, demand extensive computational resources and training time. They often require tedious hyperparameter tuning and sometimes multiple re-runs to converge to good solutions. On the other hand, we know that in supervised cases (model stitching), there exists an affine mapping between the latent spaces, suggesting the linear case is realizable – the question is how to find it efficiently. mini-vec2vec achieves that, and it has several appealing properties: it requires significantly less data than vec2vec (60k vs. 1 million samples), runs on a CPU rather than requiring extensive GPU resources, and provides more stable training dynamics.
Our method resembles the approach described in hoshen-wolf-2018-non for word embeddings, proceeding in three main stages: approximate matching, transformation fitting, and iterative refinement. This paradigm is very powerful, and we adopt a similar approach. Unfortunately, text embeddings are somewhat more complicated than word embeddings, partly because there is no small exhaustive set of objects (words) that need to be matched. Moreover, the very concept of non-overlapping datasets is somewhat contrived for word embeddings, precisely because the set of possible objects is closed and relatively small. Therefore, word embedding alignment is rarely considered in the non-overlapping datasets setup (even in multilingual matching, words mostly have a ground-truth translation). We find that applying the algorithm from hoshen-wolf-2018-non naively to our use case doesn’t work. The approximate matching stage, in particular, requires more delicate treatment in our case. The iterative refinement step is also improved by a slightly more subtle approach.
@@ 4.1 Motivation
S_A = similarity_function(C_A, C_A)
S_B = similarity_function(C_B, C_B)
@@
R_A_list.append(similarity_function(X_A, C_A))
R_B_list.append(similarity_function(X_B, C_B_permuted))
@@
We will now describe our algorithm in detail. Pseudo-code is provided in Listing LABEL:lst:full_algo.
@@ 4.2 Preprocessing
where $\boldsymbol{\mu}_{A}$ is the mean of space $A$, $\mathbf{1}_{n_{A}}$ is the vector of ones of length $n_{A}$, and $\|\cdot\|_{2}$ denotes the L2 norm.
@@ 4.3 Approximate Matching
To establish structural correspondences, we employ an anchor discovery procedure based on multiple runs of clustering and matching. We will first describe a single run, and then explain how we robustify it by ensembling multiple runs:
@@
The reason we choose to align cluster centroids is that they roughly represent the central landmarks in the dataset. Since the data comes from a shared distribution $\mathcal{D}$, we expect clusters to be similar overall, and all that remains to do is to find the alignment between them. Of course, clusters are not very consistent even with exactly the same data, so the approach needs some tolerance to errors (hence the robustification step below).
We perform k-means clustering on both embedding spaces independently to obtain cluster centroids $\{\mathbf{c}^{A}_{j}\}$ and $\{\mathbf{c}^{B}_{j}\}$ where $j=1,\ldots,k$. We then compute similarity matrices between the centroids independently in each embedding space:
@@
We then find the optimal matching between cluster centroids from different embedding spaces by solving the Quadratic Assignment Problem (QAP):
$\pi^{*}=\arg\max_{\pi\in\Pi_{k}}\;\sum_{i=1}^{k}\sum_{j=1}^{k}S^{A}_{ij}\,S^{B}_{\pi(i)\,\pi(j)}$
where $\pi:\{1,2,...,k\}\rightarrow\{1,2,...,k\}$ is a permutation from the set $\Pi_{k}$ of all permutations on $k$ elements. Due to the small number of cluster centroids, this takes only a few seconds.
Relative Representations
@@
$\mathbf{r}^{A}_{i}=[\cos(\mathbf{x}^{A}_{i},\mathbf{c}^{A}_{1}),\cos(\mathbf{x}^{A}_{i},\mathbf{c}^{A}_{2}),\ldots,\cos(\mathbf{x}^{A}_{i},\mathbf{c}^{A}_{k})]$
@@
$\mathbf{r}^{A}_{i}=[\mathbf{r}^{A}_{i,1};\mathbf{r}^{A}_{i,2};\ldots;\mathbf{r}^{A}_{i,s}]\in\mathbb{R}^{sk},\qquad\text{similarly for space }B$
This concatenation creates richer structural signatures that are more robust to noise that comes from bad alignments.
4.4 Mapping Estimation

### v4

@@ 1 Introduction
Text embeddings, also called sentence embeddings, represent chunks of text in a high-dimensional vector space. They capture complex linguistic relationships and enable downstream tasks, such as semantic similarity and information retrieval. Naturally, different embedding models induce incompatible embedding spaces that cannot be compared directly, due to differences in architecture or initialization. The output of one model looks like gibberish to another. This can be easily solved with supervised training – the translation between embedding models can be learned from embedding pairs representing the same sentences. However, the unsupervised problem, where we don’t have access to parallel embeddings of the same sentences, is much harder.
The unsupervised alignment of word embeddings has been studied extensively for a long time, often for the practical goal of aligning monolingual word embeddings from different languages. However, text embeddings have been mostly ignored. With the increasing popularity of databases of text embeddings, leakage of sensitive vector data becomes a real possibility. It has been shown that embedding inversion methods can extract the texts from the vectors alone (Song and Raghunathan, 2020; Morris et al., 2023), but they require access to the model. Unsupervised alignment can be used to extend the attack to unknown proprietary models. The attacker just translates the embeddings to a known space and applies inversion methods on the translated embeddings – extracting the text from the data (Jha et al., 2025).
@@
Recently, Jha et al. (2025) have proposed vec2vec, a method based on CycleGAN (Zhu et al., 2017) for fully unsupervised text embedding alignment. As the logic behind their method, they cite a principle sometimes called Universal Geometry or the Platonic Representation Hypothesis (Huh et al., 2024), which posits that well-trained representations, regardless of architectures or initialization, tend to converge toward geometrically similar spaces. This means that there is a way to rotate one embedding space to make it compatible with the other, and vice versa. While vec2vec is successful in finding a good alignment, the use of adversarial training is computationally intensive, requires careful hyperparameter tuning, and is subject to training instabilities like mode collapse, oscillatory behavior, or failure to converge (Saxena and Cao, 2023; Goodfellow, 2017). Moreover, adversarial methods typically demand substantial computational resources, including GPU acceleration and large amounts of training data – factors that may prohibit their utilization by most researchers and users.
We hypothesize that the transformation between the embedding spaces is of a simple form (linear, specifically), and that it can be learned without adversarial methods. Indeed, in word embedding alignment much of the work has focused on this setup as well (Artetxe et al., 2018; Hoshen and Wolf, 2018; Grave et al., 2019, inter alia). To achieve a similar goal in sentence embeddings, mini-vec2vec seeks to identify signals of similarities in the geometric structure of sentence embedding spaces. We propose a very simple pipeline (see Figure 1 for a schematic overview), which consists of three steps:
  • Approximate Matching: We match, in an unsupervised way, a small set of artificial pairs.
  • Initial Transformation: We find a linear map between the pairs and apply it to the full data.
  • Iterative Refinement: We then iteratively apply the operator, match each embedding to the closest embeddings in the other space, and learn a new operator on top of this matching.
This pipeline resembles some methods from unsupervised word embedding alignment – but text embeddings are somewhat different from word embeddings, so they require other tricks to make it work. Perhaps the most significant difference is the Approximate Matching method we rely on. At a high level, we need to find a new way to identify stable central landmarks and match them between sentence embedding spaces. Methods in word alignment implicitly rely on word vocabularies, which have special features (e.g., common words between languages, word frequencies, distributional properties of words, etc.). On the other hand, sentence embeddings can be seen as sparse samples from an “infinite” vocabulary, rendering them closer to, for instance, image embeddings.
After we have found an approximate match, the differences between word and sentence embeddings blur, and similar methods can be applied. We have developed the pipeline unaware of existing literature, so it is possible that our design decisions presented here are suboptimal or less common. Having said that, even in the Initial Alignment and Iterative Refinement steps, the design caters to specific characteristics uncommon in word embeddings, or at least less salient. For instance, when matching embeddings to their nearest neighbors, we average multiple nearest neighbors because there is no one-to-one matching between data points. This is often not the case in word alignment, since most languages, even if they are very distant, rely on closed vocabularies that significantly overlap.
@@
We find mini-vec2vec offers several key advantages: computational efficiency (very fast and runs on a commodity CPU), training stability (good performance and robust to hyperparameters), interpretability (each step has a clear geometric interpretation), and sample efficiency (60k vs 2 million samples). Through comprehensive experiments, we demonstrate that this simple approach not only matches but often exceeds the performance of the more complex adversarial method while requiring orders of magnitude less computational resources.
Figure 1: A simplified illustration of the mini-vec2vec pipeline.
@@ 2 Preliminaries
The hypothesis that neural networks converge to geometrically similar representations has been explored across multiple domains and scales. Early work by Olah (2015) has revealed striking structural similarities in representations of different convolutional networks, suggesting that the geometric organization of learned features might be more universal than previously assumed. The Platonic Representation Hypothesis, formally articulated by Huh et al. (2024), represents the most comprehensive theoretical framework for understanding Universal Geometry. The hypothesis posits that well-trained representations converge toward a shared geometric structure that reflects the underlying structure of the data domain. This means that distances (and similarities) between objects should be roughly independent of the embedding model. For example, to compute the semantic similarity between “I love your dog” and “I love my cat”, we would use cosine similarity between their embeddings. According to Universal Geometry, the answer should (theoretically) have little dependence on the model used. This makes sense intuitively, as semantic similarity originates from semantic meaning. It is a fundamentally linguistic concept learned from the data itself.
@@
Moschella et al. (2022) have operationalized the observations of Olah (2015) and proposed to use relative representations as a mathematically principled framework for comparing latent spaces. While absolute coordinates vary arbitrarily across models, we know now that relative information, like similarity between two objects, is (approximately) agnostic to the choice of embedding model. They propose to choose an agreed-upon set of data points, called anchors, that other data points will be compared to, and use them like landmarks on a map. In order to specify the location of a point, we specify how close it is to every landmark. If we have enough anchors and they are sufficiently diverse, we can “geo-locate” the point by its relative position alone. Since it requires a small set of paired anchors, the method cannot operate as-is in a fully unsupervised manner.
Denote the set of possible data points by $\mathcal{Z}$ (e.g., sentences, images, or any other modality; note these are abstract objects and not their vector representations) and the embedding function by $E:\mathcal{Z}\rightarrow\mathbb{R}^{d}$. Given a set of anchor data points $A=\{\mathbf{a}_{i}\}_{i=1}^{k}\subseteq\mathcal{Z}$, the relative representation of $\mathbf{z}\in\mathcal{Z}$ under anchors $A$ and embedding $E$ is the vector:
$\mathbf{r}_{A,E}(\mathbf{z})=[\,\operatorname{sim}(E(\mathbf{z)},E(\mathbf{a}_{1})),\,\operatorname{sim}(E(\mathbf{z}),E(\mathbf{a}_{2})),\,\ldots,\,\operatorname{sim}(E(\mathbf{z}),E(\mathbf{a}_{k}))\,],$
where $\operatorname{sim}(\cdot,\cdot)$ is a similarity function (typically cosine similarity). For two embedding models satisfying the Universal Geometry assumption, relative representations of the same data point are roughly the same, i.e., $\mathbf{r}_{A,E_{1}}(\mathbf{z})\approx\mathbf{r}_{A,E_{2}}(\mathbf{z})$.
@@ 3 Methodology
Problem Formulation
Following the setup laid out in vec2vec, let $\mathcal{X}_{A}=\{\mathbf{x}^{A}_{i}\}_{i=1}^{n_{A}}$ and $\mathcal{X}_{B}=\{\mathbf{x}^{B}_{j}\}_{j=1}^{n_{B}}$ denote two sets of $d$-dimensional embeddings. The sentences that created these embeddings are drawn from a shared underlying distribution $\mathcal{D}$, but all sentences are different and there is no overlap between the two datasets. We organize the embeddings into matrices $\mathbf{X}_{A}\in\mathbb{R}^{n_{A}\times d}$ and $\mathbf{X}_{B}\in\mathbb{R}^{n_{B}\times d}$, where each row corresponds to a single embedding vector. Our goal is to learn a transformation function $f:\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that $f(E_{A}(\mathbf{z}))\approx E_{B}(\mathbf{z})$ where $\mathbf{z}$ represents a data point $\mathbf{z}\in\mathcal{Z}$. More generally, the setup described in Jha et al. (2025) allows for learning a pair of functions $f_{A},f_{B}$ sending embeddings to a third space, such that $f_{A}(E_{A}(\mathbf{z}))\approx f_{B}(E_{B}(\mathbf{z}))$, but we will not take advantage of this flexibility here.
Motivation
In this work, we restrict our attention to linear transformations. Further, we constrain the search space to orthogonal transformations, i.e., matrices $\mathbf{W}$ that satisfy $\mathbf{W}\mathbf{W}^{T}=\mathbf{W}^{T}\mathbf{W}=\mathbf{I}$. This is motivated by the Universal Geometry hypothesis – we expect distances and similarities to be preserved by the transformation. In practice, we later relax this assumption and use an exponentially-smoothed average of orthogonal matrices to enforce orthogonality softly.
mini-vec2vec resembles many approaches in unsupervised word embedding alignment. It proceeds in three main stages: approximate matching, transformation fitting, and iterative refinement. Unfortunately, text embeddings are somewhat more complicated than word embeddings, partly because they are non-overlapping – there is no small exhaustive set (a vocabulary) of objects that need to be matched. The very concept of non-overlapping datasets is somewhat contrived for word embeddings, precisely because the set of possible objects is closed and relatively small. Even in multilingual matching, words mostly have a ground-truth one-to-one match.
Beyond the overlap between the datasets, most works (Artetxe et al., 2016, 2017, 2018; Hoshen and Wolf, 2018) rely on the statistics of words or other word-specific features (for finding a seed dictionary). Optimal transport-based methods (Alvarez-Melis and Jaakkola, 2018; Grave et al., 2019) rely heavily on matching points between the sets (often one-to-one, but this can be alleviated), and they become computationally intensive. Theoretically speaking, in a non-overlapping dataset, OT could find sentences that are (only) similar and match them to one another, and then a parametric mapping could be learned on top of them (a methodology similar to Grave et al. (2019), only applied to non-overlapping data). In practice, when Jha et al. (2025) have compared their method to multiple OT methods, in an oracle setup, they didn’t work better than the most naive baseline, indicating that even in the oracle setup, and even as a first step for learning a mapping, they are insufficient.
We will now proceed to describe our algorithm in detail (pseudo-code is provided in Listing LABEL:lst:full_algo).
@@
S_A = pairwise_similarities(C_A, C_A)
S_B = pairwise_similarities(C_B, C_B)
@@
R_A_list.append(pairwise_similarities(X_A, C_A))
R_B_list.append(pairwise_similarities(X_B, C_B_permuted))
@@ 3.1 Preprocessing
where $\boldsymbol{\mu}_{A}$ is the mean of space $A$, $\mathbf{1}_{n_{A}}$ is the vector of ones of length $n_{A}$, and $\|\cdot\|_{2}$ is the L2 norm.
@@ 3.2 Approximate Matching
To find an approximate matching, a small seed of pairs of (approximate) matches, we need to find stable landmarks that are likely to exist in both spaces, despite the non-overlapping data. The natural candidates are clusters. We conjecture that clusters represent recurring themes between the datasets and, therefore, should be consistent between the datasets. But even if we find the same set of clusters in both spaces exactly, we still need to match each cluster (centroid) to its corresponding cluster (centroid) in the other space. To match the clusters, we rely on pairwise similarities, conjecturing they encode enough information to identify the matching uniquely (this intuition is shared with Gromov-Wasserstein (Mémoli, 2011)). We match the clusters based on similarity matrices, with Quadratic Assignment Problem (QAP) solvers. These assumptions are not trivial, but the method works relatively well as-is, though often requiring a few attempts due to the stochasticity and sensitivity of the clustering and matching algorithms. To mitigate this, we run cluster-and-match multiple times and ensemble the results. We will now describe a single run, and then explain how we robustify it by ensembling.
@@
We perform k-means clustering in each embedding space independently and obtain cluster centroids $\{\mathbf{c}^{A}_{j}\}$ and $\{\mathbf{c}^{B}_{j}\}$ where $j=1,\ldots,C$. We then compute pairwise similarities in each space:
@@
We then find the optimal matching between the cluster centroids by solving the Quadratic Assignment Problem (QAP), which finds a permutation that aligns the similarity matrices optimally:
$\pi^{*}=\arg\max_{\pi\in\Pi_{C}}\;\sum_{i=1}^{C}\sum_{j=1}^{C}S^{A}_{ij}\,S^{B}_{\pi(i)\,\pi(j)}$
where $\pi:\{1,2,...,C\}\rightarrow\{1,2,...,C\}$ is a permutation from the set $\Pi_{C}$ of all permutations on $C$ elements. Due to the small number of cluster centroids, this takes only a few seconds. Conceptually, we can think of QAP in terms of cosine similarity, where we want to find the permutation that will maximize the cosine similarity between the (flattened) similarity matrices.
Option 1: Relative Representations
@@
$\mathbf{r}^{A}_{i}=[\cos(\mathbf{x}^{A}_{i},\mathbf{c}^{A}_{1}),\cos(\mathbf{x}^{A}_{i},\mathbf{c}^{A}_{2}),\ldots,\cos(\mathbf{x}^{A}_{i},\mathbf{c}^{A}_{C})]$
@@
Option 2: Alternative Approach
Curiously, we have found that it is possible to sidestep relative representations. Instead, it is possible to use the aligned centroids directly.
$\mathbf{r}^{A}_{i}=[\mathbf{c}^{A}_{1},\mathbf{c}^{A}_{2},\ldots,\mathbf{c}^{A}_{C}];\qquad\mathbf{r}^{B}_{i}=[\mathbf{c}^{B}_{\pi(1)},\mathbf{c}^{B}_{\pi(2)},\ldots,\mathbf{c}^{B}_{\pi(C)}]$
While the relative representations method learns from information derived from the entire data, the centroid-only approach requires just the centroids themselves. Both methods are efficient, and this simplified method is only slightly more efficient. It seems to do worse after the Initial Transformation step (Section 3.3), but it converges normally in the Iterative Refinement steps (Section 3.4). We emphasize that we haven’t tested it as thoroughly as the relative representation approach. For future reference, we call this variant mini-vec2vec centroid-only”. Note that the centroid-only method is also much closer in spirit to many of the alignment methods discussed above that find a seed dictionary matching (the centroids here) and build on that. In this paper, we will not focus on this variant and discuss it here only for completeness.
@@
$\mathbf{r}^{A}_{i}=[\mathbf{r}^{A}_{i,1};\mathbf{r}^{A}_{i,2};\ldots;\mathbf{r}^{A}_{i,s}]\in\mathbb{R}^{sC},\qquad\text{similarly for space }B$
This concatenation creates richer structural signatures that are more robust to the noise that comes from bad alignments.
3.3 Initial Transformation